OpenStudy (anonymous):

In the 131st running of the Kentucky Derby in 2005, there were 20 contenders. 6. How many different ways can horses place first, second, and third? WORK PLEASE!

OpenStudy (anonymous):

For this question, you're going to need to do a permutation, since order matters for first, second, and third. 20P3 = 20! / (20 - 3)! = 20! / 17! = 20 * 19 * 18 = 6,840 ways.